With explanation please.

Math 151 Test 2 Review Test 2 - Sections 3.1-3.8,4.1-4.2 Tuesday, April 3rd 1. Write the following in a+ bi form. a. (-3+5i)-(-9-8i) b. (3+5i)(-9-8i)
Answer #1

a)

(-3+5i) -(-9-8i)

= -3 +5i +9 +8i

= (-3 +9) +(5+8)i

= 6 +13i

b)

(-3+5i)(-9-8i)

= 27 +24i -45i -40i²

Use i²= -1

= 27 -21i -40(-1)

=27 -21i +40

=67 -21 i